OpenStudy (anonymous):

Simplify the squareroot of 8/9

OpenStudy (johnweldon1993):

Hint \[\large \sqrt{\frac{8}{9}} = \frac{\sqrt{8}}{\sqrt{9}}\] what do we know? well we know the square root of 9 is 3 correct? so right away we can write \[\large \frac{\sqrt{8}}{3}\] anything else we can do?

OpenStudy (anonymous):

No that's it. Thanks

OpenStudy (johnweldon1993):

Not quite :P Hint...how can we multiply 2 numbers to make 8? \[\large \frac{\sqrt{2 \times 4}}{3}\] see anything?

OpenStudy (anonymous):

Oh haha, well the sqaureroot of 4 is 2..

OpenStudy (johnweldon1993):

Indeed So the complete simplification is \[\large \frac{2\sqrt{2}}{3}\]

OpenStudy (anonymous):

So what if it was negative?

OpenStudy (johnweldon1993):

Give me an example?

OpenStudy (anonymous):

Squareroot of -9/5

OpenStudy (johnweldon1993):

Very important is it \[\large \sqrt{\frac{-9}{5}}\] or \[\large -\sqrt{\frac{9}{5}}\]

OpenStudy (anonymous):

1st one

OpenStudy (johnweldon1993):

Question, have you covered imaginary numbers?

OpenStudy (anonymous):

Kind of...but it wasn't detailed.

OpenStudy (johnweldon1993):

Alright, well as you already know you cannot take the square root of a negative number... why? well...what number squared = a negative number? There isnt one...we need to introduce imaginary numbers

OpenStudy (johnweldon1993):

So what I like to do...is ignore the fact that we have an '-' there...until the end...you'll see why :) so lets just solve \[\large \sqrt{\frac{9}{5}}\] how can you simplify that? :)

OpenStudy (anonymous):

Well the top would be three, but how would you simplify the bottom?

OpenStudy (johnweldon1993):

You wouldnt :) the square root of 5 cannot be simplified so we just have \[\large \frac{3}{\sqrt{5}}\] but remember...we ignored the '-' before so what do we do now...I'll explain :)

OpenStudy (johnweldon1993):

Let me take 1 step back to the beginning \[\large \frac{\sqrt{-9}}{\sqrt{5}}\] remember how we broke (in the first problem) the square root of 8? by saying 2 times 4? well here...think of this as \[\large \frac{\sqrt{9 \times -1}}{\sqrt{5}}\] So first....we found that the square root of 9 is 3 right? \[\large \frac{3\sqrt{-1}}{\sqrt{5}}\] now, imaginary numbers....you will see that \[\large \sqrt{-1} = i\] so our final answer is \[\large \frac{3i}{\sqrt{5}}\]

OpenStudy (johnweldon1993):

the only reason I like to ignore it until the end, is because I know all I have to do is bring along the i at the end

OpenStudy (anonymous):

Oh, well I thought in order for radicals to be simplified all the way, there shouldn't be a radical in the denominator...? Last question (:

OpenStudy (johnweldon1993):

Yes, I suppose we should :) sooooo...how would you bring the square root out of the denominator?

OpenStudy (anonymous):

I honestly don't know

OpenStudy (johnweldon1993):

Recognize the phrase "rationalize the denominator?"

OpenStudy (anonymous):

Is it where the denominator only contains rational numbers?

OpenStudy (johnweldon1993):

Indeed... and heres a hint What is \[\large \sqrt{a} \times \sqrt{a}\]?

OpenStudy (anonymous):

That #?

OpenStudy (johnweldon1993):

Indeed!! so here we have \[\large \frac{3i}{\sqrt{5}}\] if we multiply the bottom by the square root of 5...we would just have 5 right? but remember whatever we do to the bottom we do to the top...so \[\large \frac{3i}{\sqrt{5}} \times \frac{\sqrt{5}}{\sqrt{5}} = \frac{3i\sqrt{5}}{5}\] did that make sense?

OpenStudy (anonymous):

YES! Thank you so much (:

OpenStudy (johnweldon1993):

Anytime :)